OpenStudy (anonymous):

Which of the following is true about the graph of f on the interval [0,pi]? f(x)=x+sin(2x)-2 a) f(x) is increasing on the interval ((1/3)pi, (2/3)pi) b) f(x) is concave up on the interval (0, (2/3)pi) c) f(x) is decreasing on the interval (0,pi) d)f(x) has a local max at the point ((1/3)pi , (1/3)pi+(1/2)sqrt(3)-2) e)f(x) has a point of inflection at (0,-2)

OpenStudy (ranga):

You need to find first and second derivative of the function and find its behavior in the interval [0,pi]. If f'(x) is positive it is increasing. If f'(x) is negative it is decreasing. If f''(x) > 0 it is concave up. if f"(0) is 0 at (0, -2) then it has a potential point of inflection at that point, etc.

OpenStudy (anonymous):

Ok so then the first derivative of f(x) is \[1+2\cos(2x)\] and the second derivative is \[-4\sin(2x)\] But when I try to solve for zero on the first derivative I dont know exactly which are the critical numbers \[1+2\cos(2x)=0\] \[\cos(2x)=\frac{ 1 }{ 2 }\]

OpenStudy (ranga):

cos(2x) = -1/2

OpenStudy (anonymous):

I know 1/2=cos(pi/3) but does x just equal pi/6?

OpenStudy (anonymous):

or my bad a negetive 1/2

OpenStudy (anonymous):

so -1/2 of cosine is 2pi/3 so what is half of that

OpenStudy (ranga):

cos(2pi/3) = -1/2 2x = 2pi/3. x = pi/3.

OpenStudy (anonymous):

OH ok that makes much more sense now

OpenStudy (anonymous):

So then the critical number of the function is just pi/3

OpenStudy (ranga):

Also cosine repeats itself and so you have to find all solutions in the interval [0, pi] for x.

OpenStudy (anonymous):

So then I have to find where in between [0,pi] where it equals zero or its undefined?

OpenStudy (anonymous):

I dont know what you mean by how it represents itself

OpenStudy (ranga):

not represents but repeats itself. cosine is a cyclic function and so there will be multiple times it will have a value of -1/2.

OpenStudy (anonymous):

Oh ok so their are more critical values than 2pi/3 correct?

OpenStudy (anonymous):

or pi/3

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Ok then so are the cirtical numbers just pi/3 and 2pi/3 or is there more? If so how should I solve for the others?

OpenStudy (ranga):

For cosine there are infinite number of points where it is -1/2. But here the domain is [0, pi] and so x = pi/3 and 2pi/3 are the two (potential) critical points. At critical points, the function has a max/min. Evaluate second derivative at those two points.

OpenStudy (anonymous):

Ok then \[-4\sin(2(\frac{ \pi }{ 3 })<0\] and \[-4\sin(2(\frac{ 2\pi }{ 3 })>0\]

OpenStudy (anonymous):

So pi/3 is a max and 2pi/3 is a min

OpenStudy (ranga):

Yes, pi/3 is a max and 2pi/3 is a min. Therefore, we can infer the function is increasing in the interval 0, pi/3 (in order to reach max at pi/3) and decreasing in the interval pi/3, 2pi/3 (to reach min at 2pi/3).

OpenStudy (anonymous):

So my awnser choices are a) f(x) is increasing on the interval ((1/3)pi, (2/3)pi) b) f(x) is concave up on the interval (0, (2/3)pi) c) f(x) is decreasing on the interval (0,pi) d)f(x) has a local max at the point ((1/3)pi , (1/3)pi+(1/2)sqrt(3)-2) e)f(x) has a point of inflection at (0,-2) So (a) is not correct because pi/3 is a max and 2pi/3 is a min so it has to decrease between those two intervals. B is not correct because pi/3 is concave down C is not correct because their are points where f(x) increases D may be correct because pi/3 is a max but I dont know how to get to that y value E im not sure of either

OpenStudy (anonymous):

So far Im guessing its D

OpenStudy (anonymous):

Because I chose E on the quiz and it was wrong

OpenStudy (ranga):

Your conclusions about a, b and c above are correct. For d, you know the max is at x = pi/3. Put x = pi/3 and evaluate f(x).

OpenStudy (anonymous):

So \[f(\frac{ \pi }{ 3 })=\frac{ \pi }{ 3 }+\sin(2(\frac{ \pi }{ 3 })-2\]

OpenStudy (anonymous):

that equals \[\frac{ \pi }{ 3 }+\frac{ \sqrt(3) }{ 2 }-2\]

OpenStudy (anonymous):

Oh ok its simpler than Im thinking it is

OpenStudy (anonymous):

Thank you so much ranga and thanks for being patient

OpenStudy (ranga):

You are welcome Matthew. You did most of it yourself. Kudos.